OpenStudy (anonymous):

WILL MEDAL AND FAN The Latin Club charted a bus to travel to and participate in a state competition. The cost for each adult to ride the bus was $25 and the cost for each student to ride the bus was $15. A total of 17 people rode the bus. The total paid for the bus was $305. How many students and adults were on the bus?

OpenStudy (texaschic101):

25a + 15c = 305 a + c = 17 this is a substitution problem. Can you finish from here or do you need more help ?

OpenStudy (mathmate):

This is a type of problem which can be solved by a system of equations, or solved mentally as in the following \(example\): Example: students: $15, adults: $25. Total fare: $295 (REMEMBER: this is an example) for 17 persons. Solution: If all students, total fare=17*15=$255. Since we collected $295, the excess $295-255=$40 divided by ($25-$15) pays for 4 adults, so 13 students and 4 adults. (= answer to EXAMPLE).

OpenStudy (anonymous):

the number of adults is a = 5.

OpenStudy (anonymous):

The number of students is s = 12.